# Pushing JWST to the extremes: search and scrutiny of bright galaxy candidates at z15-30

### v1

@@ 4 A closer scrutiny of the selected candidates
In this section we analyse in more detail the selected candidates to assess their reliability and evaluate the possibility that their peculiar colours are instead indicative of rare lower redshift interloper populations.
@@ 4.1 The photometric redshift probability distribution
Briefly, we have run zphot adopting a set of synthetic models drawn from the BC03 library with a range of metallicities from to , , and a ?delayed? () star–formation history (see Santini et al. 2023, and M24). Nebular emission is self–consistently included following Schaerer & de Barros (2009) (see also Castellano et al. 2014), based on the template luminosity at ionizing frequencies. We have analysed our candidates with EAzY in two different ways as described in M24, i.e. 1) using only its standard set of semi-empirical templates, and 2) including also the recent set of templates from Larson et al. (2022) which are designed explicitly to reproduce the colors of high redshift galaxies. The Bagpipes runs exploit BPASS v. 2.2.1 stellar models with an upper-mass cutoff of the IMF of 300 M⊙ (Stanway & Eldridge 2018), and nebular emission computed self-consistently with CLOUDY (Ferland et al. 2013) as described by Carnall et al. (2018). Following Gandolfi et al. (2025), we increased the number of live points (i.e., the walkers used by Bagpipes in the Markov Chain Monte Carlo sampling) from the default 400 to 2000 to enhance sensitivity to strong line emitter solutions, and allowed the ionization parameter to reach log . We assume three different star-formation histories: delayed, double power law, and exponential SFH. Finally, we run CIGALE in two configurations. In the first case, we assume a SFH with a delayed component of age between 100 Myr and the age of the universe at each redshift, plus a constant burst of 10 Myr duration. The fraction of stellar mass formed in the recent burst is allowed to vary between 0 and 50% of the total assembled mass. We use BC03 templates including nebular emission, metallicity 0.02, 0.2, 1 Z⊙, and V-band extinction 05. The second CIGALE configuration exploits star-formation plus AGN templates where the stellar component is parametrized as described above and the AGN emission is based on the dale2014 module (Dale et al. 2014). The AGN fraction (fAGN), defined as the ratio of AGN luminosity to the total AGN and dust luminosities, is set as a free parameter. To summarise, from the combination of codes and assumptions adopted we have obtained eight different for all our candidates. The of the nine objects in our reference sample of F200W dropout candidates are shown in the upper panels of Fig. 3. We show in Fig. 4 the average for the extended samples of F200W- and F277W-dropout candidates.
@@
First, it is clear that all objects exhibit a solution, consistent with the color selection criteria adopted. The inferred redshifts tend to be reasonably similar among the codes, as the main spectral feature determining the high-redshift solution is the Lyman-break (coupled with the shape of the star–forming continuum immediately redward) that is essentially common in all recipes. However, all our candidates also show a lower-redshift solution, in all the runs analysed here. The low–redshift solutions are typically peaked at , suggesting that the strong observed break in the F200W band can also be ascribed to a break around the Balmer break/4000 rest-frame region, as we shall describe better in the following. In several cases z2 solutions are also viable, and generally preferred by BAGPIPES. Detailed inspection shows that this is generally due to a peculiar combination of strong emission lines that may conspire to reproduce the observed colours. We remark that, given the resulting , none of our candidates would pass a selection criterion such as 4 between different redshift solutions that has been often adopted to build luminosity functions at (e.g., Finkelstein et al. 2023; Harikane et al. 2023). Unfortunately, our candidates are simply too faint, and their photometry is built on a too small number of photometric bands with solid detections, to be unambiguously selected in the same manner. It is also clear that, while they overall provide a consistent picture of ?double peaked? solutions, there are significant differences between the adopted codes, both in terms of the breadth of the low-redshift solution and of the relative weight between the low- and high–redshift peaks. These differences arise from the different libraries adopted and probably from slight differences in the fitting procedure. This picture is confirmed by the average of the extended samples of F200W- and F277W-dropout candidates (Fig. 4).
We finally note that also depends crucially on the details of the adopted photometry. Clearly, but somewhat counterintuitively, the additional information provided by the HST photometry decreases the constraining power of . Because of the faintness of our candidates, in fact, the HST upper limits are easily satisfied by a wide class of solutions at low redshifts, eventually adding little contribution to the global while increasing the number of degrees of freedom and hence increasing the probability at low redshifts.
@@ 4.2 The nature of potential interlopers
The photometric redshift distributions described above help us unveil the physical properties of the galaxies that may contaminate our selection criteria. Adopting for simplicity the results of the zphot code, we have inspected the physical properties corresponding to the models that populate the low redshift peaks in the . The best-fit SEDs at of the nine F200W dropout candidates are shown in Fig. 3.
In the case of the two candidates in CEERS, the low-redshift best fit solution is a ?quiescent? galaxy with low specific star–formation rate and large stellar age. The remaining sources, instead, are best reproduced by star–forming models with a very high specific star–formation rate and large dust attenuation, whose red continuum and strong emission lines yield colors compatible with our selection criteria.
These two kinds of solutions are representative of the general properties of galaxy templates that typically yield low-redshift solutions for our objects, as shown in Fig. 5. We compared the solutions that provide an acceptable fit with probability for the nine F200W dropout candidates to the locus of objects in the same redshift range from the JADES-GS field. The statistically acceptable models populate regions that have a small overlap with the ones occupied by the bulk of sources in the same redshift range. Consistently with the SEDs shown in Fig. 3, the templates cover a region in the versus plane that connects quiescent, low dust models (lower left corner) with highly star-forming, dusty (upper right corner) ones. Instead, the general distribution of galaxies in this plane shows that most of the objects tend to populate the region of intermediate and . The vs plane shows that acceptable solutions include z0-4 templates with a higher dust-extinction than ?typical? sources in the same redshift range, and, in particular, are consistent with potential contamination of the F200W dropout selection by very low-mass, dusty galaxies, as previously noted by Bisigello et al. (2023) and Gandolfi et al. (2025).
@@
We remark, however, that these potential low-redshift solutions correspond to objects that would be extremely interesting to investigate. These models have stellar masses in the range , sometimes even as low as . As long as quiescent galaxies are concerned, only objects with stellar mass above have been confirmed at (Carnall et al. 2024; Glazebrook et al. 2024; Weibel et al. 2024; Pérez-González et al. 2024b). The available estimates of the stellar mass function of quiescent galaxies do not extend to these low masses, especially at these redshifts (Santini et al. 2021). These sources may also host the low-mass SMBH that are fundamental to constrain early AGN-galaxy co-evolution (Pacucci et al. 2023).
@@
4.3 The confirmed interloper UDS_56824 at =6.56
@@
One of the ten objects originally selected, the F200W-dropout candidate object UDS_56824 at R.A.=34.454893 deg., Dec=-5.215586 deg (see Fig. 2), has been recently observed with the NIRSpec PRISM in the framework of the CAPERS program. The observation has been carried out by adopting a NRSIRS2 readout pattern, standard 3-shutter “slits”, and a 3-point nodding for a total integration time of 5690 secs. The data were reduced with the STScI Calibration Pipeline222https://jwst-pipeline.readthedocs.io/en/latest/index.html version 1.13.4 as described in detail in Arrabal Haro et al. (2023a) (see also Castellano et al. 2024; Napolitano et al. 2025). Although the spectrum covers only the region at 3m, the redshift can be accurately measured to be =6.560.01 from the weighted average centroid of the H, H, [O III] lines which are clearly detected (Fig. 6). We measured line fluxes with a Gaussian fit after linearly extrapolating the continuum emission at the line position (see Napolitano et al. 2024, 2025, for details) finding FHα=16.2 2.1 10, FHβ=2.1 0.9 10, F[OIII]4959=2.1 1.1 10, and F[OIII]5007=1.26 0.18 10.
The Balmer ratio implies a high dust attenuation with =0.850.46 when assuming a Calzetti et al. (2000) attenuation law. Its [O III]/H=6.02.7 puts the object in the star-forming region of the mass-excitation diagram (Juneau et al. 2014), although at the border with the AGN locus, such that a contribution from a dust-obscured active nucleus cannot be excluded.
@@
We checked the ALMA archive finding that the position of UDS_56824 has been observed by project #2015.1.01074.S (PI H. Inami) that we analyse as follows. We start from the calibrated measurement set and use the CASA tclean function to create a continuum map using the four 2 GHz-wide spectral windows, with an effective central frequency of 343 GHz (i.e., 870 m). We apply natural weighting to the visibilities and test the effect of adding a uvtaper value to artificially increase the beam size. A 3 detection is found at the NIRCam position of UDS_56824, with an integrated flux density of mJy (see Fig. 6). This galaxy is fainter than all SCUBA-2-selected galaxies in UDS (Geach et al. 2017) and all previous ALMA detections in the field reported in Dudzevičiūtė et al. (2020), suggesting that a significant population of dusty galaxies at z6 may have been missed by previous submillimeter surveys.
@@
The combination of spectroscopic and photometric information demonstrates that UDS_56824 is an example of a highly attenuated high-redshift starburst galaxy. Its large UV attenuation 4.5 mag, makes it a significant outlier in the relation (McLure et al. 2018), similarly to CEERS-93316 (Arrabal Haro et al. 2023b) and CEERS-14821 (Bisigello et al. 2025). In fact, UDS_56824 shows that even at z6.5, some galaxies may be dominated by dust-obscured star formation (note that the uncorrected H-based SFR is 2-4 M⊙ yr-1). The other known interloper CEERS-93316 at z=4.9 is remarkably similar to UDS_56824 in terms of both extinction (0.6 mag) and SFR (30 M⊙ yr-1), but with a lower stellar mass log8.4-9 (Arrabal Haro et al. 2023b).
@@
These findings are consistent with our analysis in Sect. 4.2 pointing toward dusty star-forming galaxies up to z7-8 as a potential source of contamination of the F200W-dropout sample. Similarly to CEERS-93316, the SED of UDS_56824 resembles that of a z15 LBG due to the combination of a red, attenuated continuum and line emission. However, it is apparent that a proper characterization of these interloper populations is extremely difficult and requires spectroscopic follow-up observations. In fact, a redshift of z6.5 is a secondary peak in our (Fig. 7) but the most likely alternative solution was a similarly dusty, star-forming object with log9 at z4 whose broad-band fluxes are boosted by a different combination of emission lines. Moreover, the properties of UDS_56824 appear extreme compared to the locus of alternative solutions of the other 15 candidates, while CEERS-93316 is more in line with the expectations (Fig. 5).
4.4 Critical assessment of the photometric sample
@@
We exploit the analysis described above to perform a detailed evaluation of the objects in our reference sample of F200W-dropouts. We first note that four objects in our sample consistently show a probability for the high-redshift solution which is significantly lower than any of the low/intermediate redshift peaks: COSMOS_31168, COSMOS_76919, COSMOS_84213 and COSMOS_118438. In the case of COSMOS_31168 and COSMOS_84213, this is likely due to a marginal detection (SNR2) in one HST band each, albeit both objects are non-detected in both the ACS and WFC3 stacks that we used for our selection. In addition, both these sources as well as COSMOS_76919 show a drop in the F410M band similarly to UDS_56824, thus making dusty star-forming solutions more likely. On the contrary, COSMOS_118438 has a slightly higher flux in F410M than in F356W and F444W, which also leads SED-fitting codes to prefer strong-line emitting templates at z6. In fact, we tested that the probability of the solution increases after removing the F410M from the fit for all the aforementioned sources. The remaining five sources in our sample have high-redshift solutions as significant as the low-redshift ones in most of the SED-fitting runs. In particular, CEERS_15937, CEERS_17384 and COSMOS_107923 show a high consistency among all codes and recipes. However, there are compelling reasons to consider their reliability with caution. The two CEERS objects appear somewhat similar to CEERS-93316, as they have a clear peak at and a red continuum in the F277W/F356W/F444W bands, leading us to suspect that they are members of the same 5 overdensity (Naidu et al. 2022b; Arrabal Haro et al. 2023b) together with other known quiescent (de Graaff et al. 2025) and dusty (Zavala et al. 2023; Bisigello et al. 2025) galaxies. Finally, COSMOS_107923, albeit being the brightest in the sample falls in a region of relatively shallow image depth and therefore has a corresponding limited sampling of the spectral break. This source also has -1.2 which is redder than the UV slopes measured in spectroscopically confirmed objects at (Roberts-Borsani et al. 2024b, a), with the only notable exception of the X-ray emitting AGN GHZ9 at z=10.145 (=-1.10.12, Napolitano et al. 2024).

### v2

@@ 4 A closer scrutiny of the selected candidates
In this section we analyse in more detail the selected candidates to assess their reliability and evaluate the possibility that their peculiar colours are instead indicative of rare lower redshift interloper populations. Although we focus here on the 5 candidates, we remark that very similar arguments can be made for the 5 sources that have been confirmed to be interlopers, as discussed in Sect. 4.3.
@@ 4.1 The photometric redshift probability distribution
Briefly, we have run zphot adopting a set of synthetic models drawn from the BC03 library with a range of metallicities from to , , and a ?delayed? () star–formation history (see Santini et al. 2023, and M24). Nebular emission is self–consistently included following Schaerer & de Barros (2009) (see also Castellano et al. 2014), based on the template luminosity at ionizing frequencies. We have analysed our candidates with EAzY in two different ways as described in M24, i.e. 1) using only its standard set of semi-empirical templates, and 2) including also the recent set of templates from Larson et al. (2023) which are designed explicitly to reproduce the colors of high redshift galaxies. The Bagpipes runs exploit BPASS v. 2.2.1 stellar models with an upper-mass cutoff of the IMF of 300 M⊙ (Stanway & Eldridge 2018), and nebular emission computed self-consistently with CLOUDY (Ferland et al. 2013) as described by Carnall et al. (2018). Following Gandolfi et al. (2025), we increased the number of live points (i.e., the walkers used by Bagpipes in the Markov Chain Monte Carlo sampling) from the default 400 to 2000 to enhance sensitivity to strong line emitter solutions, and allowed the ionization parameter to reach log . We assume three different star-formation histories: delayed, double power law, and exponential SFH. Finally, we run CIGALE in two configurations. In the first case, we assume a SFH with a delayed component of age between 100 Myr and the age of the universe at each redshift, plus a constant burst of 10 Myr duration. The fraction of stellar mass formed in the recent burst is allowed to vary between 0 and 50% of the total assembled mass. We use BC03 templates including nebular emission, metallicity 0.02, 0.2, 1 Z⊙, and V-band extinction 05. The second CIGALE configuration exploits star-formation plus AGN templates where the stellar component is parametrized as described above and the AGN emission is based on the dale2014 module (Dale et al. 2014). The AGN fraction (fAGN), defined as the ratio of AGN luminosity to the total AGN and dust luminosities, is set as a free parameter. To summarise, from the combination of codes and assumptions adopted we have obtained eight different for all our candidates. The of the five objects in our reference sample of F200W dropout candidates are shown in the upper panels of Fig. 3. We show in Fig. 4 the average for the extended samples of F200W- and F277W-dropout candidates.
@@
First, it is clear that all objects exhibit a solution, consistent with the color selection criteria adopted. The inferred redshifts tend to be reasonably similar among the codes, as the main spectral feature determining the high-redshift solution is the Lyman-break (coupled with the shape of the star–forming continuum immediately redward) that is essentially common in all recipes. However, all our candidates also show a lower-redshift solution, in all the runs analysed here. The low–redshift solutions are typically peaked at , suggesting that the strong observed break in the F200W band can also be ascribed to a break around the Balmer break/4000 rest-frame region, as we shall describe better in the following. In several cases z2 solutions are also viable, and generally preferred by BAGPIPES. Detailed inspection shows that this is generally due to a peculiar combination of strong emission lines that may conspire to reproduce the observed colours. We remark that, given the resulting , none of our candidates would pass a selection criterion such as 4 between different redshift solutions that has been often adopted to build luminosity functions at (e.g., Finkelstein et al. 2024a; Harikane et al. 2023). Unfortunately, our candidates are simply too faint, and their photometry is built on a too small number of photometric bands with solid detections, to be unambiguously selected in the same manner. It is also clear that, while they overall provide a consistent picture of ?double peaked? solutions, there are significant differences between the adopted codes, both in terms of the breadth of the low-redshift solution and of the relative weight between the low- and high–redshift peaks. These differences arise from the different libraries adopted and probably from slight differences in the fitting procedure. This picture is confirmed by the average of the extended samples of F200W- and F277W-dropout candidates (Fig. 4).
We note that also depends crucially on the details of the adopted photometry. Clearly, but somewhat counterintuitively, the additional information provided by the HST photometry decreases the constraining power of . Because of the faintness of our candidates, in fact, the HST upper limits are easily satisfied by a wide class of solutions at low redshifts, eventually adding little contribution to the global while increasing the number of degrees of freedom and hence increasing the probability at low redshifts. These results are based on the ASTRODEEP-JWST catalogs, which provide homogeneous photometric measurements for all sources. We find that the MIRI constraints discussed in Sect. 3.2 do not significantly change the scenario described above. The maintain the same double-peaked nature with both high- and low-redshift solutions when including the MIRI upper limits, and, in the case of CEERS_17384, the tentative F770W detection. In fact, the MIRI non-detections are poorly constraining for all our sources, with the exception of COSMOS_107923 for which the limit in the F770W increases the probability of passive solutions with respect to dusty star-forming ones in the low-redshift peak. In fact, the MIRI imaging should be 1-2 mag deeper in the F770W and F1000W bands, and 3-4 mag deeper at 15m to yield meaningful constraints for the other candidates. Finally, when including additional NIRCam medium-band photometry for JADES-GN_9538 we find an overall similar but with a slightly increased probability at z13-15 due to the marginal detection in the F182M band.
@@ 4.2 The nature of potential interlopers
The photometric redshift distributions described above help us unveil the physical properties of the galaxies that may contaminate our selection criteria. Adopting for simplicity the results of the zphot code, we have inspected the physical properties corresponding to the models that populate the low redshift peaks in the . The best-fit SEDs at of the F200W dropout candidates are shown in Fig. 3.
In the case of the candidate in CEERS, the low-redshift best fit solution is a ?quiescent? galaxy with low specific star–formation rate and large stellar age. The remaining sources, instead, are best reproduced by star–forming models with a very high specific star–formation rate and large dust attenuation, whose red continuum and strong emission lines yield colors compatible with our selection criteria.
These two kinds of solutions are representative of the general properties of galaxy templates that typically yield low-redshift solutions for our objects, as shown in Fig. 5. We compared the solutions that provide an acceptable fit with probability for the F200W-dropout candidates to the locus of objects in the same redshift range from the JADES-GS field. The statistically acceptable models populate regions that have a small overlap with the ones occupied by the bulk of sources in the same redshift range. Consistently with the SEDs shown in Fig. 3, the templates cover a region in the versus plane that connects quiescent, low dust models (lower left corner) with highly star-forming, dusty (upper right corner) ones. Instead, the general distribution of galaxies in this plane shows that most of the objects tend to populate the region of intermediate and . The vs plane shows that acceptable solutions include z0-4 templates with a higher dust-extinction than ?typical? sources in the same redshift range, and, in particular, are consistent with potential contamination of the F200W dropout selection by very low-mass, dusty galaxies, as previously noted by Bisigello et al. (2023) and Gandolfi et al. (2025).
@@
We remark, however, that these potential low-redshift solutions correspond to objects that would be extremely interesting to investigate. These models have stellar masses in the range , sometimes even as low as . As long as quiescent galaxies are concerned, only objects with stellar mass above have been confirmed at (Carnall et al. 2024; Glazebrook et al. 2024; Weibel et al. 2025a; Pérez-González et al. 2025a). The available estimates of the stellar mass function of quiescent galaxies do not extend to these low masses, especially at these redshifts (Santini et al. 2021). These sources may also host the low-mass SMBH that are fundamental to constrain early AGN-galaxy co-evolution (Pacucci et al. 2023).
@@
4.3 The confirmed interlopers
Five of the ten objects originally selected as F200W-dropouts have been recently observed with the NIRSpec PRISM in the framework of the CAPERS program, and found to be interlopers at z2-7 (empty symbols in Fig. 2). We describe in the following their physical properties as inferred from the combination of NIRSpec spectroscopy (Fig. 6) and NIRCam photometry (Sect. B). These observations confirm that contaminants of our selection criteria are found in the populations presented in Sect. 4.2, namely dusty star—forming or quiescent galaxies with low stellar mass.
@@ 4.3 The confirmed interlopers
The spectroscopic observations have been carried out by adopting a NRSIRS2 readout pattern, standard 3-shutter “slits”, and a 3-point nodding. The data were reduced with the STScI Calibration Pipeline333https://jwst-pipeline.readthedocs.io/en/latest/index.html version 1.13.4 as described in detail in Arrabal Haro et al. (2023a) (see also Castellano et al. 2024; Napolitano et al. 2025). We determine the spectroscopic redshift from the weighted average centroid of the detected emission lines, and measured line fluxes with a Gaussian fit after linearly extrapolating the continuum emission at the line position (see Napolitano et al. 2024, 2025, for details).
4.3.1 COSMOS_31168 at =2.29
The spectrum of COSMOS_31168 (R.A.= 150.1809331 deg., Dec=2.2607514 deg.) shows one clear emission line at 1.64m, a lower significance emission feature at 2.15m and a faint continuum. The most plausible redshift solution is =2.294, interpreting the aforementioned emission lines as the unresolved [O III] doublet (with flux 9.71.9 10) and H (=3.10.5 10). The H line is not detected, with a 2 flux upper limit of 0.75 10. The corresponding limit on the Balmer ratio results in =0.32. We performed an SED-fitting of the object with zphot by fixing the redshift at the spectroscopic value and adopting the aforementioned constraint on , finding that the object is fitted as a low-mass (log = 8), low SFR (=0.07 M⊙ yr-1) dusty object.
4.3.2 COSMOS_35731 at =3.66
The CAPERS spectrum of this source (R.A.= 150.133574 deg., Dec=2.2710158 deg.) has prominent H (28.60.2 10), [O III] (8.70.2 10) and H (4.80.2 10) emission, yielding a robust spectroscopic redshift =3.657. The Balmer decrement implies a high extinction (0.60.04), with a corrected SFR measured from the H luminosity of 1.9 0.1 M⊙ yr-1. The SED-fitting performed at fixed redshift and constraining within the 2 range inferred from spectroscopy yields a stellar mass of only log = 7.6.
4.3.3 CEERS_15937 at =4.91
The spectrum of CEERS_15937 (R.A.= 214.9442826 deg., Dec= 52.8358477 deg.) shows continuum emission with a clear break at 2.15m. The shape of the continuum and the faint emission detectable down to 1m shows that the object is not at very high-redshift. Due to the lack of evident emission features, we fitted the spectrum with BAGPIPES assuming a double power-law SFH, and leaving the redshift as a free parameter. We find that the object is fitted by a low-mass (log = 8.9 0.05) galaxy at =4.91, with low extinction (0.02) and negligible SFR, corresponding to an upper limit of sSFR-11.4 Gyr-1.
4.3.4 COSMOS_76919 at =5.18
This object (R.A.= 150.1845568 deg., Dec= 2.3535079 deg.) is constrained to be at =5.182 from the detection of the H (2.20.5 10), [O III] (4.80.3 10), and H (0.4 0.2 10) lines. Similarly to COSMOS_31168 and COSMOS_35731, the object turns out to be a dusty (0.50.3), low-mass (log = 8.2 0.4), and low SFR (=0.3 0.1 M⊙ yr-1) object on the basis of the Balmer decrement, H luminosity and constrined SED-fitting performed with zphot.
4.3.5 UDS_56824 at =6.56
@@
The spectrum of object UDS_56824 (R.A.=34.454893 deg., Dec=-5.215586 deg) covers only the region at 3m, with a total integration time of 5690 secs. The redshift can be accurately measured to be =6.560.01 from the H, H, [O III] lines, which are clearly detected with fluxes FHα=16.2 2.1 10, FHβ=2.1 0.9 10, F[OIII]4959=2.1 1.1 10, and F[OIII]5007=1.26 0.18 10. The Balmer ratio implies a high dust attenuation with =0.850.46. Its [O III]/H=6.02.7 puts the object in the star-forming region of the mass-excitation diagram (Juneau et al. 2014), although at the border with the AGN locus, such that a contribution from a dust-obscured active nucleus cannot be excluded. We checked the ALMA archive finding that the position of UDS_56824 has been observed by project #2015.1.01074.S (PI H. Inami) that we analyse as follows. We start from the calibrated measurement set and use the CASA tclean function to create a continuum map using the four 2 GHz-wide spectral windows, with an effective central frequency of 343 GHz (i.e., 870 m). We apply natural weighting to the visibilities and test the effect of adding a uvtaper value to artificially increase the beam size. A 3 detection is found at the NIRCam position of UDS_56824, with an integrated flux density of mJy (see Fig. 6).
@@
4.4 Critical assessment
@@
These findings are consistent with our analysis in Sect. 4.2 indicating dusty star-forming galaxies and low-mass passive objects up to z7-8 as a potential source of contamination of the F200W-dropout sample. However, it is apparent that a proper characterization of these interloper populations is extremely difficult and requires spectroscopic follow-up observations. In fact, CEERS_15937 is the only case for which the yielded an alternative solution in good agreement with the spectroscopic value (Fig. 16). Instead, the redshifts of the other four objects do not match the primary, low-redshift peaks in our , although the most likely alternative solutions are similarly dusty, star-forming objects whose broad-band fluxes are boosted by different combinations of emission lines.
Similarly to CEERS-93316 (Arrabal Haro et al. 2023b), the SEDs of COSMOS_31168, COSMOS_35731, COSMOS_76919 and UDS_56824 resemble that of z15 LBGs due to the combination of a red, attenuated continuum and line emission. Instead, CEERS_15937 enters the F200W-droput selection due to the well-known ambiguity between Lyman and Balmer break detection through broad-band photometry. Their position in the vs. , and in the vs. planes are in line with the expectations based on the (Fig. 5). The four dusty star-forming interlopers are significant outliers in the relation (McLure et al. 2018), similarly to CEERS-93316 and CEERS-14821 (Bisigello et al. 2025). Object UDS_56824 appears to be an extreme case in terms of redshift and stellar mass compared to the locus of alternative solutions of the 15 candidates. The other known interlopers at z2-5 are similar to UDS_56824 in terms of extinction and sSFR, but with a lower stellar mass log9. In fact, UDS_56824 shows that even at z6.5, some galaxies may be dominated by dust-obscured star formation (note that the uncorrected H-based SFR is 2-4 M⊙ yr-1). Finally, this ALMA-detected galaxy is fainter than all SCUBA-2-selected sources in UDS (Geach et al. 2017) and all previous ALMA detections in the field reported in Dudzevičiūtė et al. (2020), suggesting that a significant population of dusty galaxies at z6 may have been missed by previous submillimeter surveys.
Object CEERS_15937 at z=4.91 is remarkable, being among the most distant confirmed passive objects. Most importantly, it is the least massive known to date, the only two comparable cases being GS-z5-Q1 and COS-z5-Q1 with a 3 times larger stellar mass (log9.5-9.6, Baker et al. 2025). Similarly to GS-z5-Q1 and COS-z5-Q1, CEERS_15937 is a member of a known 5 overdensity (Naidu et al. 2022b; Arrabal Haro et al. 2023b) together with other known quiescent (de Graaff et al. 2025) and dusty (Zavala et al. 2023; Bisigello et al. 2025) galaxies. The number density and formation pathway of low-mass quiescent galaxies at high-redshift are presently unknown (e.g., Merlin et al. 2025). An in-depth investigation of CEERS_15937 and similar objects may shed light on the role of enviroment at early times and on their connection with the recently discovered class of ?mini-quenched? galaxies at high-redshift (Strait et al. 2023; Looser et al. 2024).
@@
We exploit the analysis described above to perform a detailed evaluation of the objects in our reference sample of F200W-dropouts. We first note that two objects in our sample consistently show a probability for the high-redshift solution which is significantly lower than any of the low/intermediate redshift peaks: COSMOS_84213 and COSMOS_118438. In the case of COSMOS_84213, this is likely due to a marginal detection (SNR2) in one HST band each, albeit it is non-detected in both the ACS and WFC3 stacks that we used for our selection. In addition, this source show a drop in the F410M band similarly to COSMOS_31168, COSMOS_76919, and UDS_56824, thus making dusty star-forming solutions more likely. On the contrary, COSMOS_118438 has a slightly higher flux in F410M than in F356W and F444W, which also leads SED-fitting codes to prefer strong-line emitting templates at z6. In fact, we tested that the probability of the solution increases after removing the F410M from the fit for all the aforementioned sources. The remaining three sources in our sample have high-redshift solutions as significant as the low-redshift ones in most of the SED-fitting runs. In particular, CEERS_17384 and COSMOS_107923 show a high consistency among all codes and recipes. However, there are compelling reasons to consider their reliability with caution. The CEERS object appear somewhat similar to CEERS_15937, as it has a clear peak corresponding to a passive solution at , leading us to suspect that it is a member of the same 5 overdensity (Naidu et al. 2022b; Arrabal Haro et al. 2023b). Finally, COSMOS_107923, albeit being the brightest in the sample falls in a region of relatively shallow image depth and therefore has a corresponding limited sampling of the spectral break. This source also has -1.2 which is redder than the UV slopes measured in spectroscopically confirmed objects at (Roberts-Borsani et al. 2024, 2025a), with the only notable exception of the X-ray emitting AGN GHZ9 at z=10.145 (=-1.10.12, Napolitano et al. 2024).